If the area of a circle is 720 cm2, what is the area of the sector if the central angle measures 12degrees?


A storage tank in the shape of a cube is to be insulated using material that costs $2.50 per square foot. If the edge of the cube is 12 feet, how much will the insulation cost?

Answer by Alan3354(69443)

If the area of a circle is 720 cm2, what is the area of the sector if the central angle measures 12degrees?
= 720*(12/360) = 24 sq cm
-----------------------
A storage tank in the shape of a cube is to be insulated using material that costs $2.50 per square foot. If the edge of the cube is 12 feet, how much will the insulation cost?
Area = 6s^2 = 6*12^2 = 864 sq ft
cost = 2.5*864 = $2160